graph the equation x-4y=2
Found 2 solutions by hkelson, Alan3354:
Answer by hkelson(7)

In order to graph this equation it must be changed into slope-intercept form. To do this you will do the following steps:
x - 4y = 2 : Subtract x from both sides (this is to get the y by itself)
-4y = -x + 2 : Divide both sides by -4
y = 1/4x - 2/4 : Reduce to lowest terms
y = 1/4x - 1/2
If you have a graphing calculator just plug it into the y= and hit the graph button. Or if you have to write it down on graphing paper find the y-intercept and place a point and then use the slope to find the next point, up 1 and over 4.

Answer by Alan3354(69443)

graph the equation x-4y=2
-----------------
I did. It's a straight line, which is to be expected from a linear equation (linear is from line).
You can graph it too, if you go to www.padowan.dk.com/graph/ and DL the (free) software.